OpenStudy (anonymous):

Sin 46 degrees equals x over 14. What is x?

OpenStudy (richyw):

\[\sin{45^\circ}=\frac{x}{14}\]

OpenStudy (richyw):

oops. read the number wrong\[\sin{46^{\circ}}=\frac{x}{14}\] can you see what to do now? you will probably want a calculator for it. if it was 45 degrees you wouldn't need one.

OpenStudy (richyw):

Alright well if you are still not getting it then just remember that \(\sin{46^{\circ}}\) is just a constant number. So you have \(\frac{x}{14}\) equals some number. But you don't want x divided by 14, you want x on its own. so you need to multiply x by the denominator, in this case multiply it by 14. An remember that whatever you to do one side, you must do to the other.

OpenStudy (richyw):

so multiply left and right hand side by 14 \[14\sin{46^{\circ}}=x\]is the answer. you can calculate that out. It should be around \(x=7\sqrt{2}\) since that is the answer for \(x=14\sin{45^{\circ}}\)

OpenStudy (anonymous):

sin46=x/14 x=14sin46 x=14*0.719 x=10.07

OpenStudy (richyw):

there you go, someone has a calculator! haha. I guess I could have used wolfram